OpenStudy (anonymous):

Root of 2 + Root of 6 +Root of 8 How do you get from that too 3root2+root of 6? I'll help you if you help me.

hartnn (hartnn):

\[\sqrt{8}=\sqrt{2*4}=\sqrt{2*2^2}=2\sqrt{2}\] now can u get that?

OpenStudy (anonymous):

yeahh, then?

hartnn (hartnn):

so, \[\sqrt{2}+\sqrt{6}+2\sqrt{2}=3\sqrt{2}+\sqrt{6}\] got this?

OpenStudy (anonymous):

okay thanks!

hartnn (hartnn):

welcome :)